# Asynchronous Feedback Network for Perceptual Point Cloud Quality Assessment

### v1

@@ IV-C Performance Comparison
Cross-Dataset Evaluation. The cross-dataset evaluation is conducted to test the generalizability of the NR-PCQA methods when encountering various data distributions. In Table III, we first train the compared models on the LS-PCQA and test the trained model on SJTU-PCQA and WPC, and then swap the training set and test set. Note that we use the whole database for training or testing. For each validation, we record the result with minimal training loss and report the results in Table III. From the table, we can see that: i) AFNet achieves good performance when trained on LS-PCQA and WPC, which demonstrates the generalizability of our model in most cases. ii) Almost all methods (including AFNet) show poor performance when trained on SJTU-PCQA. This is largely due to the limited and extremely biased content in SJTU-PCQA (5 out of 9 contents are about human body).
IV-D Ablation Studies
Ablation Study for the Key Modules. To verify the effectiveness of the key modules in the AFNet, we further investigate the contribution of different components and report the results in Table IV. In the table, the index (1) and (2) denote using only global branch or local branch for quality prediction respectively. (3) indicates one dual-branch network without feedback connection (i.e., replacing the DRConv in Fig. 4 (b) with $3\times 3$ convolution), while (4) represents the full model. From the table, we have the following observations: i) Seeing (1), (2), and (3), the dual-branch structure performs significantly better than the single-branch structure, demonstrating the complementarity of the two features; ii) Seeing (3) and (4), the feedback module further brings a noticeable gain, which verifies the effectiveness of the feedback mechanism in visual quality assessment.
@@ IV-D Ablation Studies
Ablation Study for the Region Count. We test the performance of AFNet under different $n$ values to investigate the influence of the division granularity and report the results in Fig. 8. We see that SJTU-PCQA, WPC and LS-PCQA respectively provide the best performance with $n=6,8,$ and $10$. This may be due to the different amount of content in the three databases. SJTU-PCQA has limited and extremely biased content, so few divisions can distinguish different semantic regions while too many divisions cause over-segmentation for the same region. In comparison, WPC and LS-PCQA have more references (20 for WPC and 85 for LS-PCQA) and richer content including different object categories, for which cases large $n$ values can benefit the model performance.
TABLE V: Ablation study for the input image types on the WPC dataset.
Index Texture Depth Occupancy PLCC SROCC RMSE
(1) 0.8815 0.8850 10.6951
(2) 0.3374 0.2128 21.3751
(3) 0.8821 0.8826 10.6614
(4) 0.3908 0.2548 20.8975
(5) 0.8971 0.8979 9.9474
(6) 0.8974 0.8987 9.9707
Ablation Study for the Kernel Size. We test the performance of AFNet under different $k$ values to investigate the influence of the kernel size of region-wise filters and report the results in Fig. 9. We see that when the kernel size is too large (e.g., $k=7$), the model performance will decrease. The reason may be that too large kernel sizes influence feature learning near to semantic region boundaries. More specifically, on the semantic region boundaries, a large kernel may convolute with pixels from different semantic regions. Threrefore, a relatively small kernel size (e.g., $k=3$) is preferred by the proposed model.
@@
Ablation Study for the Input Image Types. Three types of image are generated during the preprocessing stage. The depth images serve as a supplement to texture images in the network input stage, while the occupancy images help perform multi-view fusion and region-aware feedback. We conduct experiments to validate the effectiveness of the three types of input and list the results in Table V. In the table, index (1) and (2) represent using only texture or depth images for quality prediction respectively. (3), (4), and (5) denote pairwise combinations of different image types, including ”texture+depth”, ”depth+occupancy”, and ”texture+occupancy”, while (6) indicates the full model. Note that the framework using only occupancy images is infeasible because they are only used to generate some weights and masks. From the table, we can observe: i) based on (1) and (2), model using texture images performs far better than that using depth images. It is reasonable because texture images generally contain visual information than depth images. ii) Comparing (1) with (3), introducing the depth images brings a marginal performance gain in terms of PLCC and RMSE. This is due to the fact that geometry distortions can also be reflected in texture images, therefore reducing the importance of depth information. iii) Seeing (1) and (5), (3) and (6), the introduction of occupancy images leads to noticeable improvements for the three criteria, demonstrating the importance of occupancy information in our model.
@@
Ablation Study for the Loss Functions. The proposed network is trained using the regression loss $L_{reg}$, the feature disentangling loss $L_{dis}$, and the quality ranking loss $L_{rank}$. We evaluate the effect of each loss function and report the results in Table VI. In the table, index (1), (2) and (3) represent using single loss term for quality prediction respectively. (3), (4), and (5) denote pairwise combinations of different loss terms, including $L_{reg}+L_{dis}$, $L_{dis}+L_{rank}$, and $L_{reg}+L_{rank}$, while (6) indicates the full loss. From the table, it is observed: i) seeing (1), (4), and (5), both $L_{dis}$ and $L_{rank}$ benefit the model performance, demonstrating the effectiveness of the two loss functions. ii) According to (6) and (7), the full loss provide marginal gain over the combination of $L_{reg}+L_{rank}$. The reason may be that $L_{rank}$ also achieves feature disentanglement to some extent, which weakens the effect of $L_{dis}$. Considering $L_{dis}$ only measure the linear similarity between global and local features, one possible avenue for further improvement is to minimize non-linear dependence between two features, such as mutual information.
@@ V Conclusion
In this paper, we aim to solve the problem of global-local feature learning in NR-PCQA and propose an asynchronous feedback network based on the human visual mechanisms. Taking into account the guiding role of global feature in quality perception, our network is implemented by a dual-branch architecture connected by a feedback module. Utilizing the attention maps generated by the global branch, we perform region-aware convolution to derive the local feature. A coarse-to-fine learning strategy is also adopted to fuse the global and local feature into the final quality prediction. The proposed AFNet shows a consistent and reliable correlation with subjective evaluation on three PCQA benchmarks. Further ablation studies have supported the model design by examining its key modules and parameter values.
TABLE VI: Ablation study for the loss functions on the WPC dataset.
Index $L_{reg}$ $L_{dis}$ $L_{rank}$ PLCC SROCC RMSE
(1) 0.8851 0.8868 10.4514
(2) 0.1985 0.1133 22.2610
(3) 0.3724 0.2317 21.1203
(4) 0.8898 0.8908 10.3274
(5) 0.4247 0.3418 20.5666
(6) 0.8970 0.8981 10.0299
(7) 0.8974 0.8987 9.9707

### v2

@@ IV-D Ablation Studies
Ablation Study for the Region Count. We test the performance of AFQ-Net under different $n$ values to investigate the influence of the division granularity and report the results in Fig. 8. We see that SJTU-PCQA, WPC and LS-PCQA respectively provide the best performance with $n=6,8,$ and $10$. This may be due to the different amount of content in the three databases. SJTU-PCQA has limited and extremely biased content, so few divisions can distinguish different semantic regions while too many divisions cause over-segmentation for the same region. In comparison, WPC and LS-PCQA have more references (20 for WPC and 85 for LS-PCQA) and richer content including different object categories, for which cases large $n$ values can benefit the model performance.
Ablation Study for the Kernel Size. We test the performance of AFQ-Net under different $k$ values to investigate the influence of the kernel size of region-wise filters and report the results in Fig. 9 (a) and (b). We see that when the kernel size is too large (e.g., $k=7$), the model performance will decrease. Especially, $7\times 7$ kernel causes inferior results on LS-PCQA. Further looking at this case, we find that larger kernel size in DRConv will generate larger element values in the local feature, which will make the local feature contribute more to the prediction, thus affecting the convergence speed of the network. More specifically, given the local feature $f_{l}$, we define its average absolute value of as $\mu_{l}=\mathrm{mean}(\mathrm{abs}(f_{l}))$.We have illustrated the SROCC (on the training set) trend and the $\mu_{l}$ trend versus the training epoch for $3\times 3$ and $7\times 7$ kernel in Fig. 10 (a) and (b). From Fig. 10 (b), we can see the $\mu_{l}$ of $7\times 7$ kernel is significantly larger than that of $3\times 3$ kernel; according to Fig. 10 (a), the corresponding SROCC on the training set converges more slowly. In fact, DRConv tends to generate element values in the range of $[0,1]$ for convolution kernels, therefore causing larger results with increasing kernel size. Meanwhile, the complexity of LS-PCQA make the network harder to converge compared to WPC and SJTU-PCQA. A reasonable solution is to add a normalization factor for different kernel size, such as $\frac{1}{9}$ for $3\times 3$ kernel and $\frac{1}{49}$ for $7\times 7$ kernel. We report the SROCC and the $\mu_{l}$ trend of two normalized kernels in Fig. 10 (c) and (d). We can see the normalized $3\times 3$ and $7\times 7$ kernels present similar trends in both two sub-figures. Meanwhile, the final [PLCC,SROCC,RMSE] criteria for the two normalized kernels are $[0.6826,0.6706,0.5858]$ and $[0.6755,0.6664,0.5962]$ respectively, which demonstrates that the normalization can effectively eliminate the performance drop caused by large kernels.
Figure 10: SROCC trend and $\mu_{l}$ trend versus the training epoch for different kernels.
@@
Ablation Study for the Input Image Types. Three types of image are generated during the preprocessing stage. The depth images serve as a supplement to texture images in the network input stage, while the occupancy images help perform multi-view fusion and region-aware feedback. We conduct experiments to validate the effectiveness of the three types of input and list the results in Table VI. In the table, index (1) and (2) represent using only texture or depth images for quality prediction respectively. (3), (4), and (5) denote pairwise combinations of different image types, including “texture+depth”, “depth+occupancy”, and “texture+occupancy”, while (6) indicates the full model. Note that the framework using only occupancy images is infeasible because they are only used to generate some weights and masks. From the table, we can observe: i) based on (1) and (2), model using texture images performs far better than that using depth images. It is reasonable because texture images generally contain visual information than depth images. ii) Comparing (1) with (3), introducing the depth images brings a marginal gain and even impairs the model performance. This is due to the fact that geometry distortions can also be reflected in texture images, therefore reducing the importance of depth information. Especially, most distortion types in LS-PCQA do not cause any geometry deformation, which result in relatively inferior performance when injecting depth information. iii) Seeing (1) and (5), (3) and (6), the introduction of occupancy images leads to noticeable improvements for the three criteria, demonstrating the importance of occupancy information in our model.
@@
Ablation Study for the Loss Functions. The proposed network is trained using the regression loss $L_{reg}$, the feature disentangling loss $L_{dis}$, and the quality ranking loss $L_{rank}$. We evaluate the effect of each loss function and report the results in Table VII. In the table, index (1), (2) and (3) represent using single loss term for quality prediction respectively. (3), (4), and (5) denote pairwise combinations of different loss terms, including $L_{reg}+L_{dis}$, $L_{dis}+L_{rank}$, and $L_{reg}+L_{rank}$, while (6) indicates the full loss. From the table, it is observed: i) seeing (1), (4), and (5), both $L_{dis}$ and $L_{rank}$ benefit the model performance, demonstrating the effectiveness of the two loss functions. ii) According to (6) and (7), the full loss provides noticeable gain over the combination of $L_{reg}+L_{rank}$ on LS-PCQA but presents marginal gain on WPC. The reason may be that $L_{rank}$ also achieves feature disentanglement to some extent, which weakens the effect of $L_{dis}$. Considering $L_{dis}$ only measures the linear similarity between global and local features, one possible avenue for further improvement is to minimize non-linear dependence between two features, such as mutual information.
TABLE VI: Ablation study for the input image types
Index Texture Depth Occupancy WPC LS-PCQA
PLCC SROCC PLCC SROCC
(1) 0.8815 0.8850 0.6734 0.6583
(2) 0.3374 0.2128 0.2354 0.0485
(3) 0.8821 0.8826 0.6655 0.6508
(4) 0.3908 0.2548 0.2398 0.0154
(5) 0.8971 0.8979 0.6907 0.6744
(6) 0.8974 0.8987 0.6902 0.6796
@@ V Conclusion
In this paper, we aim to solve the problem of global-local feature learning in NR-PCQA and propose an asynchronous feedback network based on the human visual mechanisms. Taking into account the guiding role of global feature in quality perception, our network is implemented by a dual-branch architecture connected by a feedback module. Utilizing the attention maps generated by the global branch, we perform region-aware convolution to derive the local feature. A coarse-to-fine learning strategy is also adopted to fuse the global and local feature into the final quality prediction. The proposed AFQ-Net shows a consistent and reliable correlation with subjective evaluation on three PCQA benchmarks. Further ablation studies have supported the model design by examining its key modules and parameter values.
TABLE VII: Ablation study for the loss functions
Index $L_{reg}$ $L_{dis}$ $L_{rank}$ WPC LS-PCQA
PLCC SROCC PLCC SROCC
(1) 0.8851 0.8868 0.6696 0.6538
(2) 0.1985 0.1133 0.3097 0.0236
(3) 0.3724 0.2317 0.2428 0.0660
(4) 0.8898 0.8908 0.6704 0.6571
(5) 0.4247 0.3418 0.2802 0.1379
(6) 0.8970 0.8981 0.6704 0.6600
(7) 0.8974 0.8987 0.6902 0.6796